An investor invested a total of $2200 in two mutual funds. One earned 5% profit while the other earned a 3% profit. If the investor's total profit was $86, how much was invested in each mutual fund?
Answer by mananth(16946)

Part I 3.00% per annum ------------- Amount invested =x
Part II 5.00% per annum ------------ Amount invested = y
2200
Interest----- 86.00

Part I 3.00% per annum ---x
Part II 5.00% per annum ---y
Total investment
x + 1 y= 2200 -------------1
Interest on both investments
3.00% x + 5.00% y= 86
Multiply by 100
3 x + 5 y= 8600.00 --------2
Multiply (1) by -3
we get
-3 x -3 y= -6600.00
Add this to (2)
0 x 2 y= 2000
divide by 2
y = 1000
Part I 3.00% $ 1200
Part II 5.00% $ 1000

CHECK
1200 --------- 3.00% ------- 36.00
1000 ------------- 5.00% ------- 50.00
Total -------------------- 86.00

m.ananth@hotmail.ca